OpenStudy (solomonzelman):

I want to be able to 720 and + the math on the SAT, can you give me a problem that I can't do in about 1 minute? (No links, please)

HanAkoSolo (jamierox4ev3r):

first of all that's great..hmm let me see.... Let the function \(f\) be defined by \(f(x)=x^2+18\). If \(m\) is a positive number such that \(f(2m)=2f(m)\). What is the value of \(m\)? ^ Solve that :P

OpenStudy (solomonzelman):

Can you explain this one to me? (Anyone?)

OpenStudy (solomonzelman):

HELP!

HanAkoSolo (jamierox4ev3r):

its an SAT question :P that's all that is given

OpenStudy (solomonzelman):

Can you help me? I forgot how to do this?

HanAkoSolo (jamierox4ev3r):

well I guess here's a problem that ya can't do in 1 minute- mission accomplished ^_^ anyhow, just use the model of x; m is just a value within the function, this problem is basically just reasoning

OpenStudy (solomonzelman):

I wasn't trying to brag, dude! I just need to be able to do all of the SAT problems....

HanAkoSolo (jamierox4ev3r):

lol nooo well here let me help you with this one. what is 18 divided by 2?

OpenStudy (solomonzelman):

Really? Lets see huuuum! 9 right? I am really not sure (...don't mock me please, this website is not for that)

HanAkoSolo (jamierox4ev3r):

don't overthink it :P no seriously, I'm not trying to! this problem has very complicated structure for sure! but the answer is pretty easy to come across to

HanAkoSolo (jamierox4ev3r):

and the square root of 9 is...

OpenStudy (solomonzelman):

I am joking, .... although you do look like u are....

HanAkoSolo (jamierox4ev3r):

sorry >.<

HanAkoSolo (jamierox4ev3r):

but \(\sqrt{9}\) is ...

OpenStudy (solomonzelman):

3 OK, dude! I said SAT "Problem" they don't ask that on the SAT! Although they could ask sq root of 9 +4, simplify, or something along these lines....

HanAkoSolo (jamierox4ev3r):

....and 3 is literally your answer!!! you just have to know what to plug in and whatnot, and after that its really simple operations

HanAkoSolo (jamierox4ev3r):

they DO ask these questions on the SAT! its problem number 18 , section 2 of the SAT. Page 891 in the official SAT guide. Need I say more? :P

OpenStudy (solomonzelman):

OK you know that I could.... can you ask something harder, and explain how to do it, if I seem to be an able to do it?

OpenStudy (solomonzelman):

unable... i meant

HanAkoSolo (jamierox4ev3r):

well that question I just asked was categorized as Super Hard in the book, its wording is meant to confuse you. That's the evils of the SAT . so uhm okay lets see.. In the figure below, the perimeter of the triangle is \(4+2\sqrt{2}\). what is the value of x? Choices: a.) 2 b.) 4 c.) \(\sqrt{2}\) d.) \(2\sqrt{2}\) e.) \(2+\sqrt{2}\)

OpenStudy (solomonzelman):

45-45-90 triangle?

OpenStudy (solomonzelman):

then 2

HanAkoSolo (jamierox4ev3r):

the brolem doesn't state it explicitly

OpenStudy (solomonzelman):



HanAkoSolo (jamierox4ev3r):

*problem* but yes indeed......good job, and good assumption lol

OpenStudy (solomonzelman):

x and x are equal and the perimiter is 4 plus 2 tims sq root of 2...

HanAkoSolo (jamierox4ev3r):

if you want i can share some more later, but right now I have to run, so seeya :P

OpenStudy (solomonzelman):

Not assumtion, reading

OpenStudy (solomonzelman):

More problems?